OpenStudy (princeevee):

Multiply and simplify: –3ab3 • (–4a4b5) A. –12a5b8 B. –12a4b15 C. 12a5b8 D. 12a4b15

OpenStudy (texaschic101):

lets take this one step at a time what is -3 * -4 ??

OpenStudy (princeevee):

-12

OpenStudy (texaschic101):

try again....a negative times a negative is a positive

OpenStudy (princeevee):

nvm 12

OpenStudy (texaschic101):

correct....so that eliminates the first 2 now... what is a * a^4 ??

OpenStudy (princeevee):

a^5

OpenStudy (texaschic101):

correct.. when multiplying exponents with the same base, you add the exponents so now we have 12a^5 now what is b^3 * b^5 ?

OpenStudy (princeevee):

b^8

OpenStudy (texaschic101):

correct...leaving us with .. 12a^5b^8 questions ?

OpenStudy (princeevee):

well 1 more

OpenStudy (texaschic101):

I can try...what ya got

OpenStudy (princeevee):

Multiply and simplify: 3xy4z2 • 5x2yx A. 8x4y5z2 B. 15x2y4z2 C. 15x4y5z2 D. 15(xyz)11

OpenStudy (texaschic101):

ok...start with just the numbers what is 3 * 5 ?

OpenStudy (princeevee):

15

OpenStudy (texaschic101):

what is x * x^2 ?

OpenStudy (princeevee):

x^3

OpenStudy (texaschic101):

ok..I am not getting any of your answer choices I am getting 15x^3y^5z^3

OpenStudy (princeevee):

there exactly how they are

OpenStudy (texaschic101):

wait..I read the problem wrong...hold on

OpenStudy (texaschic101):

3 * 5 = 15 x * x^2 * x = x^4 y * y^4 = y^5 and don't forget the z^2 leaving us with : 15x^4y^5z^2

OpenStudy (princeevee):

thx!

OpenStudy (texaschic101):

sure..sorry about the confusion :)